OpenStudy (1018):

I need a refresher with my integration, help me with this problem:

OpenStudy (1018):

\[\int\limits_{}^{}cosx^2 xdx\]

OpenStudy (welshfella):

use substitution put u = x^2 du / dx = 2x

OpenStudy (1018):

guys im sorry! it;s cos^2 xdx

OpenStudy (1018):

really really sorry

OpenStudy (welshfella):

for this one use the identity cos^2 x = (1/2) ( cos 2x + 1)

OpenStudy (welshfella):